# Improving Human Verification of LLM Reasoning through Interactive Explanation Interfaces

### v2

@@ Abstract
The reasoning capabilities of Large Language Models (LLMs) have led to their increasing employment in several critical applications, particularly education, where they support problem-solving, tutoring, and personalized study. Chain-of-thought (CoT) reasoning capabilities [1, 2] are well-known to help LLMs decompose a problem into steps and explore the solution spaces more effectively, leading to impressive performance on mathematical and reasoning benchmarks. As the length of CoT tokens per question increases substantially to even thousands of tokens per question [1], it is unknown how users could comprehend LLM reasoning and detect errors or hallucinations. To address this problem and understand how reasoning can improve human-AI interaction, we present three new interactive reasoning interfaces: interactive CoT (iCoT), interactive Program-of-Thought (iPoT), and interactive Graph (iGraph). That is, we ask LLMs themselves to generate an interactive web interface wrapped around the original CoT content, which may be presented in text (iCoT), graphs (iGraph) or code (iPoT). This interface allows users to interact with and provide a novel experience in reading and validating the reasoning chains of LLMs. Across a study of 125 participants, interactive interfaces significantly improve user performance. Specifically, iGraph users score the highest error detection rate ($85.6\%$), followed by iPoT ($82.5\%$), iCoT ($80.6\%$), all outperforming standard CoT ($73.5\%$). Interactive interfaces also lead to faster user validation time—iGraph users are faster ($57.9$ secs per question) than the users of iCoT and iPoT ($60$ secs) and the standard CoT ($64.7$ secs). A post-study questionnaire shows that users prefer iGraph, citing its superior ability to enable them to follow the LLM’s reasoning. We discuss the implications of these results and provide recommendations for the future design of reasoning models. Try our interactive interfaces on HuggingFace.
@@ 1 Introduction
Large language models (LLMs) are now woven into high-stakes, real-world workflows, including education, where they assist with problem solving, tutoring, and personalized study [3, 4, 5]. A key driver of their recent success on complex tasks is chain-of-thought (CoT) prompting [6]: eliciting step-by-step intermediate reasoning often yields dramatic accuracy gains on arithmetic, commonsense, and symbolic reasoning benchmarks, with state-of-the-art performance on GSM8K and related math datasets. However, the current format for presenting LLM explanations may be far from optimal [7].
@@
Figure 1: Given a GSM8K question, LLMs typically provide step-by-step reasoning followed by the final answer. However, such output presentation is often static and long, posing a higher cognitive load to users and leading to slower and more erroneous answer verification. In contrast, we prompt LLMs to generate an interactive HTML/JavaScript application wrapped around the reasoning. This interface enables users to verify the reasoning more efficiently via tools of (a) navigation buttons (inspired by those in common IDEs) and (b) colored highlights [8].
@@
While CoT outputs contain rich information, their long-text presentation introduces significant challenges for human interpretation [9, 10, 11]. In practice, especially in modern reasoning LLMs [1], CoT often appears as a long “wall of text” that can be time-consuming, obscure structure, increase cognitive load, and make error detection harder precisely when it matters. For example, when an explanation contains a subtle mistake in one intermediate step, detecting that error often requires carefully parsing multiple sentences and mentally reconstructing the reasoning flow [12]. More broadly, the explainable AI (XAI) literature shows that explanation techniques are frequently designed around what researchers can produce rather than what end users need; users instead report wanting practically useful information that improves collaboration, trust calibration, and feedback loops with AI systems [13, 14, 15]. This gap motivates rethinking how reasoning is surfaced to end-users in contexts like education, where clarity, verification, and engagement are essential.
@@
Prior work has demonstrated that the design of LLM explanations, which includes factors such as visual structure, interactivity, and information hierarchy, can significantly influence users’ trust and understanding [16, 17, 18, 19]. However, most existing research on CoT focuses on improving the reasoning process within the model rather than optimizing the presentation of the reasoning [20, 21]. As a result, LLM explanations remain largely static, text-heavy, and difficult to navigate [22]. This disconnect raises critical questions: How can explanation formats be redesigned to reduce cognitive load and enable users to effectively detect errors and understand LLM-generated reasoning steps?
@@ 2 Related work
LLMs in Education with Chain-of-Thought (CoT) Prompting. A key prompting technique for LLMs is CoT prompting, which elicits the LLM’s reasoning steps and improves its performance on multi-step tasks [23]. A recent study shows that few-shot CoT dramatically boosts arithmetic, commonsense, and symbolic reasoning accuracy [2, 3]. In educational settings, however, how explanations are presented matters. Despite clear opportunities for tutoring and feedback, Khosravi et al. [24] highlights persistent concerns about transparency, interpretability, and over-reliance in LLMs. Beyond transparency, CoT reasoning tends to be quite lengthy, which can increase cognitive load and slow down the user’s comprehension [25]. Given the concerns about the interpretability and cognitive demands of LLM-generated reasoning, researchers have begun to explore how reasoning quality and user understanding can be evaluated more systematically. A growing body of work has adopted verification accuracy, which indicates the users’ ability to correctly judge whether a model’s reasoning is valid, as a key indicator of comprehension in LLM explanations [26, 27, 28, 29, 19]. Further, a recent study argues that explanations are only useful when they enable users to verify predictions step-by-step, rather than simply accepting plausible reasoning at face value [30]. In our study, we aim to address this question by measuring how effectively people can detect reasoning errors across different explanation formats, allowing us to assess not only the quality of the reasoning itself but also how various interface designs influence users’ ability to verify and understand it.
@@
Interactivity in LLM Explanations. Recent studies have emphasized the importance of interactivity in AI explanations to improve deeper user engagement and understanding[31, 32, 33]Teso et al. [34] provides a comprehensive overview of interactive machine learning, which shows that explanations are most effective when users can query, edit, or refine them rather than passively receive them. They argue that interaction transforms explanations from static outputs into dynamic learning tools that encourage users’ cognitive engagement. Similarly, another work reviews human-centered explainability and finds that interactive explanation interfaces significantly enhance engagement, comprehension, and retention compared to one-way text outputs [35]. Their analysis shows that many existing AI systems remain static and fail to leverage interactive elements that foster user engagement and a deeper understanding of the AI reasoning process. In educational settings, a study demonstrates that interactive LLM-driven question–answer systems increase learners’ focus and cognitive engagement without introducing cognitive load [36]. Together, these studies highlight that interactivity in AI explanations not only improves the user’s understanding but also keeps the users engaged.
@@ 3.1 Designing Alternative Response Formats for LLM Reasoning
The goal of this project is to design and evaluate alternative response formats for LLM-generated reasoning for math problems. While LLMs have demonstrated strong capabilities in producing step-by-step reasoning through CoT prompting, prior research shows that these outputs often are lengthy, text-heavy passages that are difficult for students to parse and verify in practice [37]. Therefore, we investigate whether alternative reasoning formats can enhance comprehension and reduce the cognitive load of users. To address these challenges, we design three alternative interactive reasoning formats that transform the same underlying content into different modes of presentation. We now introduce our three alternative explanation formats alongside the baseline explanation format.
Traditional Chain-of-Thought (CoT) serves as the baseline explanation format. It presents reasoning in a purely text-based form, where the model expresses its thought process step-by-step in natural language. Each step builds upon the previous one and forms a continuous logic flow that leads to the final solution[2]. The interface includes three main sections: the problem statement, the model’s step-by-step explanation, and the final answer (see Fig. 2 A). This format represents the standard way large language models naturally generate explanations without additional structural guidance or interactivity.
@@
Interactive Program-of-Thought (iPoT) shows explanations in a structured, code-like format(see Fig. 2 C). Each reasoning step is represented as a line of pseudo-code, where variables are defined, updated, and computed systematically. This format allows users to trace how each intermediate value is derived through reasoning steps. Playback controls enable users to execute the reasoning “step-by-step,” similar to debugging or running a script. This interface design draws inspiration from the “Program of Thoughts Prompting” framework proposed by Chen et al. [38], which demonstrated that representing reasoning as executable programs can improve interpretability and consistency in LLMs.
@@ 3.2 Dataset Construction
To evaluate how alternative explanation formats aid users in detecting LLM-generated errors, we construct a dataset derived from GSM8K, a widely used benchmark of grade school math word problems [39]. GSM8K is chosen because it contains a large and diverse set of multi-step reasoning tasks that resemble problems students might encounter in classroom and tutoring settings. Each GSM8K problem typically involves 3–8 steps of reasoning with natural language explanations, which makes it well-suited for transformation into interactive formats.
A central aspect of our design is the controlled injection of errors into the explanations. Rather than relying on naturally occurring model mistakes, following Kamoi et al. [40], we adopt a taxonomy of nine error categories: Calculation (CA), Counting (CO), Context Value (CV), Contradictory Step (CS), Missing Step (MS), Hallucination (HA), Unit Conversion (UC), Operator (OP), and Formula Confusion (FC). By systematically injecting each of these error types, we creat a consistent and fine-grained framework that enable us to assess not only whether participants could recognize when an explanation is incorrect, but also whether certain categories of errors are more easily detected across different interface formats. For each error type, we randomly sample 50 problems from GSM8K. We ensure that each selected problem contains between four to seven reasoning steps. This range is chosen because shorter problems may make error detection trivial, while longer problems risk overburdening participants. Altogether, this process produced 450 erroneous explanations (9 error types × 50 examples each). To balance the dataset and establish a baseline for comparison, we also sample an additional 50 correct explanations from GSM8K. Like the erroneous cases, these problems also contain 4–7 reasoning steps in length and served to evaluate whether participants could correctly recognize valid reasoning when presented in different formats. By combining erroneous and correct cases, we obtain a dataset of 500 explanations in total.
@@ 3.5 Data Flow and Interface Generation
Our proposed pipeline from problem input to rendered explanation involves two key stages, as illustrated in Fig.3. First, each math problem and its solution are passed to the LLM (Claude 3.7 Sonnet [41] in all our experiments), which produces a tagged intermediate representation. This stage ensures that the reasoning content is captured in a structured format. In the second stage, the tagged representation is combined with the interface-specific templates, where the tags provide the logical content and the templates enforce layout rules and structural consistency. This combined input is then passed to another LLM, which generated the final rendered interface in one of four formats: traditional CoT, iCoT, iPoT, or iGraph.
@@ 4.2 Evaluation Metrics
We collect quantitative and qualitative measures to capture both performance outcomes and user experiences: i) Verification Accuracy [19]: For each explanation, we record whether participants correctly labeled the solution as correct or incorrect, providing a direct indicator of how effectively each interface supports error detection and overall comprehension; ii) Error Localization [40]: In cases where the user identifies an explanation as incorrect, they are required to indicate the specific step at which the error occurred. This allows us to assess not only error detection but also participants’ ability to pinpoint the exact reasoning breakdown (avoiding the user answering by guessing); iii) Response Time [8]: The system automatically logs the total time spent on each explanation, which we use as a proxy for cognitive effort, with longer times potentially indicating greater difficulty in comprehension or in identifying error; iv) Interaction Behavior [42]: All interface interactions, including play, pause, step forward, and step back, are recorded at the click level. These logs gave us insights into how participants navigate the explanations; and v) Subjective Feedback [43]: After completing the tasks, participants rate their assigned interface on dimensions such as clarity, ease of use, and overall satisfaction using Likert-scale questions. They are also invited to provide open-ended feedback about what they found helpful or confusing, helping us understand participants’ perceptions of strengths and weaknesses in each format.
@@ 4.3 Post survey Questionnaire
After the experiment, participants will complete 2 questionnaires displayed in Appendix B to assess both their interaction experience and their perceptions of the explanation quality. The questionnaires are designed to address the study’s research questions regarding interpretability, efficiency, and engagement across different explanation formats. All items use a 7-point Likert scale (1 = Strongly Disagree, 7 = Strongly Agree) to capture nuanced user perceptions. Two types of questionnaires are given: Design Feature and a Post-Study. The question sets are inspired by established measures of engagement, usability, and interpretability in human-AI interaction research and are adapted for the design features of our interfaces [44, 43, 45].
@@ 5.1 Quantitative Result
Results in Fig. 5 demonstrate that interactive explanation formats substantially improve participants’ ability to verify reasoning accuracy compared to the traditional CoT. In particular, the iGraph interface achieves the highest verification accuracy ($85.6\%$), followed by the iPoT ($82.5\%$) and iCoT ($80.6\%$) formats, whereas CoT only achieves $73.5\%$. This pattern indicates that interactive explanation formats help users better detect and localize logical faults. These findings align with recent work emphasizing the importance of human-centered explainability, which explains understanding as a process of active answer-seeking rather than passive reception [46].
Next, we use a Kruskal–Wallis ANOVA to examine whether participants’ ability to identify the specific step containing an error varies across the four explanation formats. The analysis reveals a statistically significant difference among formats ($H(3)=13.12,p=0.004$). Pairwise Mann–Whitney U tests with Bonferroni correction show that iCoT($U=300.8,p=0.045$), iPoT ($U=271.5,p=0.021$), and iGraph ($U=153.0,p<0.001$) interfaces all achieve significantly higher wrong-step identification accuracy compared to the traditional CoT format. Results in Fig. 6 show that IGraph has achieved the highest wrong-step identification accuracy ($85.2\%$), followed by iPoT ($80.1\%$) and iCoT ($79.3\%$), compared to only $66.1\%$ reached by traditional CoT. The higher wrong-step identification accuracy on iGraph further suggests that interactive explanations helped users understand each step more thoroughly. Prior works in explainable artificial intelligence (XAI) argue that unstructured explanations often overwhelm users with text and often make it hard for users to keep track of each reasoning step [47]. Our findings support these arguments by showing that interactive explanations allow users to accurately identify the origin of an error, rather than simply noticing that one exists. Further, visualizing reasoning steps as interconnected nodes and edges transforms purely textual information into a two-dimensional spatial format, making the logic flow easier to follow and understand (this is further verified by our post survey rating results in Fig. 8).

### v3

@@ Abstract
The reasoning capabilities of Large Language Models (LLMs) have led to their increasing employment in several critical applications, particularly education, where they support problem-solving, tutoring, and personalized study. Chain-of-thought (CoT) reasoning capabilities [guo2025deepseek, wei2022chain] are well-known to help LLMs decompose a problem into steps and explore the solution spaces more effectively, leading to impressive performance on mathematical and reasoning benchmarks. As the length of CoT tokens per question increases substantially to even thousands of tokens per question [guo2025deepseek], it is unknown how users could comprehend LLM reasoning and detect errors or hallucinations. To address this problem and understand how reasoning can improve human-AI interaction, we present three new interactive reasoning interfaces: interactive CoT (iCoT), interactive Program-of-Thought (iPoT), and interactive Graph (iGraph). That is, we ask LLMs themselves to generate an interactive web interface wrapped around the original CoT content, which may be presented in text (iCoT), graphs (iGraph) or code (iPoT). This interface allows users to interact with and provide a novel experience in reading and validating the reasoning chains of LLMs. Across a study of 125 participants, interactive interfaces significantly improve user performance. Specifically, iGraph users score the highest error detection rate ($85.6\%$), followed by iPoT ($82.5\%$), iCoT ($80.6\%$), all outperforming standard CoT ($73.5\%$). Interactive interfaces also lead to faster user validation time—iGraph users are faster ($57.9$ secs per question) than the users of iCoT and iPoT ($60$ secs) and the standard CoT ($64.7$ secs). A post-study questionnaire shows that users prefer iGraph, citing its superior ability to enable them to follow the LLM’s reasoning. We discuss the implications of these results and provide recommendations for the future design of reasoning models. Try our interactive interfaces on HuggingFace.
@@ 1 Introduction
Large language models (LLMs) are now woven into high-stakes, real-world workflows, including education, where they assist with problem solving, tutoring, and personalized study [kojima2022large, wang2024large, forootani2025survey]. A key driver of their recent success on complex tasks is chain-of-thought (CoT) prompting [wang-etal-2023-towards]: eliciting step-by-step intermediate reasoning often yields dramatic accuracy gains on arithmetic, commonsense, and symbolic reasoning benchmarks, with state-of-the-art performance on GSM8K and related math datasets. However, the current format for presenting LLM explanations may be far from optimal [agarwal2024faithfulness].
@@
Figure 1: Given a GSM8K question, LLMs typically provide step-by-step reasoning followed by the final answer. However, such output presentation is often static and long, posing a higher cognitive load to users and leading to slower and more erroneous answer verification. In contrast, we prompt LLMs to generate an interactive HTML/JavaScript application wrapped around the reasoning. This interface enables users to verify the reasoning more efficiently via tools of (a) navigation buttons (inspired by those in common IDEs) and (b) colored highlights [nguyen2025hot].
@@
While CoT outputs contain rich information, their long-text presentation introduces significant challenges for human interpretation [kunz2024properties, hedlin2025got, turpin2023language]. In practice, especially in modern reasoning LLMs [guo2025deepseek], CoT often appears as a long “wall of text” that can be time-consuming, obscure structure, increase cognitive load, and make error detection harder precisely when it matters. For example, when an explanation contains a subtle mistake in one intermediate step, detecting that error often requires carefully parsing multiple sentences and mentally reconstructing the reasoning flow [mishra2025investigating]. More broadly, the explainable AI (XAI) literature shows that explanation techniques are frequently designed around what researchers can produce rather than what end users need; users instead report wanting practically useful information that improves collaboration, trust calibration, and feedback loops with AI systems [miller2019explanation, abdul2018trends, kaur2020interpreting]. This gap motivates rethinking how reasoning is surfaced to end-users in contexts like education, where clarity, verification, and engagement are essential.
@@
Prior work has demonstrated that the design of LLM explanations, which includes factors such as visual structure, interactivity, and information hierarchy, can significantly influence users’ trust and understanding [yang2020visual, alicioglu2022survey, jiang2023graphologue, nguyen2024interpretable]. However, most existing research on CoT focuses on improving the reasoning process within the model rather than optimizing the presentation of the reasoning [wang2024chain, wang2022towards]. As a result, LLM explanations remain largely static, text-heavy, and difficult to navigate [zhao2024explainability]. This disconnect raises critical questions: How can explanation formats be redesigned to reduce cognitive load and enable users to effectively detect errors and understand LLM-generated reasoning steps?
@@ 2 Related work
LLMs in Education with Chain-of-Thought (CoT) Prompting. A key prompting technique for LLMs is CoT prompting, which elicits the LLM’s reasoning steps and improves its performance on multi-step tasks [zhang2022automatic]. A recent study shows that few-shot CoT dramatically boosts arithmetic, commonsense, and symbolic reasoning accuracy [wei2022chain, kojima2022large]. In educational settings, however, how explanations are presented matters. Despite clear opportunities for tutoring and feedback, khosravi2022explainable highlights persistent concerns about transparency, interpretability, and over-reliance in LLMs. Beyond transparency, CoT reasoning tends to be quite lengthy, which can increase cognitive load and slow down the user’s comprehension [nayab2024concise]. Given the concerns about the interpretability and cognitive demands of LLM-generated reasoning, researchers have begun to explore how reasoning quality and user understanding can be evaluated more systematically. A growing body of work has adopted verification accuracy, which indicates the users’ ability to correctly judge whether a model’s reasoning is valid, as a key indicator of comprehension in LLM explanations [narayanan2018humans, linder2021level, nourani2020don, lage2019human, nguyen2024interpretable]. Further, a recent study argues that explanations are only useful when they enable users to verify predictions step-by-step, rather than simply accepting plausible reasoning at face value [fok2024search]. In our study, we aim to address this question by measuring how effectively people can detect reasoning errors across different explanation formats, allowing us to assess not only the quality of the reasoning itself but also how various interface designs influence users’ ability to verify and understand it.
@@
Interactivity in LLM Explanations. Recent studies have emphasized the importance of interactivity in AI explanations to improve deeper user engagement and understanding[kim2020improved, guesmi2024interactive, nguyen2024allowing]teso2023leveraging provides a comprehensive overview of interactive machine learning, which shows that explanations are most effective when users can query, edit, or refine them rather than passively receive them. They argue that interaction transforms explanations from static outputs into dynamic learning tools that encourage users’ cognitive engagement. Similarly, another work reviews human-centered explainability and finds that interactive explanation interfaces significantly enhance engagement, comprehension, and retention compared to one-way text outputs [zhang2025human]. Their analysis shows that many existing AI systems remain static and fail to leverage interactive elements that foster user engagement and a deeper understanding of the AI reasoning process. In educational settings, a study demonstrates that interactive LLM-driven question–answer systems increase learners’ focus and cognitive engagement without introducing cognitive load [ozdel2025examining]. Together, these studies highlight that interactivity in AI explanations not only improves the user’s understanding but also keeps the users engaged.
@@ 3.1 Designing Alternative Response Formats for LLM Reasoning
The goal of this project is to design and evaluate alternative response formats for LLM-generated reasoning for math problems. While LLMs have demonstrated strong capabilities in producing step-by-step reasoning through CoT prompting, prior research shows that these outputs often are lengthy, text-heavy passages that are difficult for students to parse and verify in practice [plevris2023chatbots]. Therefore, we investigate whether alternative reasoning formats can enhance comprehension and reduce the cognitive load of users. To address these challenges, we design three alternative interactive reasoning formats that transform the same underlying content into different modes of presentation. We now introduce our three alternative explanation formats alongside the baseline explanation format.
Traditional Chain-of-Thought (CoT) serves as the baseline explanation format. It presents reasoning in a purely text-based form, where the model expresses its thought process step-by-step in natural language. Each step builds upon the previous one and forms a continuous logic flow that leads to the final solution[wei2022chain]. The interface includes three main sections: the problem statement, the model’s step-by-step explanation, and the final answer (see Fig. 2 A). This format represents the standard way large language models naturally generate explanations without additional structural guidance or interactivity.
@@
Interactive Program-of-Thought (iPoT) shows explanations in a structured, code-like format(see Fig. 2 C). Each reasoning step is represented as a line of pseudo-code, where variables are defined, updated, and computed systematically. This format allows users to trace how each intermediate value is derived through reasoning steps. Playback controls enable users to execute the reasoning “step-by-step,” similar to debugging or running a script. This interface design draws inspiration from the “Program of Thoughts Prompting” framework proposed by chen2022program, which demonstrated that representing reasoning as executable programs can improve interpretability and consistency in LLMs.
@@ 3.2 Dataset Construction
To evaluate how alternative explanation formats aid users in detecting LLM-generated errors, we construct a dataset derived from GSM8K, a widely used benchmark of grade school math word problems [cobbe2021training]. GSM8K is chosen because it contains a large and diverse set of multi-step reasoning tasks that resemble problems students might encounter in classroom and tutoring settings. Each GSM8K problem typically involves 3–8 steps of reasoning with natural language explanations, which makes it well-suited for transformation into interactive formats.
A central aspect of our design is the controlled injection of errors into the explanations. Rather than relying on naturally occurring model mistakes, following kamoi2024evaluating, we adopt a taxonomy of nine error categories: Calculation (CA), Counting (CO), Context Value (CV), Contradictory Step (CS), Missing Step (MS), Hallucination (HA), Unit Conversion (UC), Operator (OP), and Formula Confusion (FC). By systematically injecting each of these error types, we creat a consistent and fine-grained framework that enable us to assess not only whether participants could recognize when an explanation is incorrect, but also whether certain categories of errors are more easily detected across different interface formats. For each error type, we randomly sample 50 problems from GSM8K. We ensure that each selected problem contains between four to seven reasoning steps. This range is chosen because shorter problems may make error detection trivial, while longer problems risk overburdening participants. Altogether, this process produced 450 erroneous explanations (9 error types × 50 examples each). To balance the dataset and establish a baseline for comparison, we also sample an additional 50 correct explanations from GSM8K. Like the erroneous cases, these problems also contain 4–7 reasoning steps in length and served to evaluate whether participants could correctly recognize valid reasoning when presented in different formats. By combining erroneous and correct cases, we obtain a dataset of 500 explanations in total.
@@ 3.5 Data Flow and Interface Generation
Our proposed pipeline from problem input to rendered explanation involves two key stages, as illustrated in Fig.3. First, each math problem and its solution are passed to the LLM (Claude 3.7 Sonnet [Anthropic] in all our experiments), which produces a tagged intermediate representation. This stage ensures that the reasoning content is captured in a structured format. In the second stage, the tagged representation is combined with the interface-specific templates, where the tags provide the logical content and the templates enforce layout rules and structural consistency. This combined input is then passed to another LLM, which generated the final rendered interface in one of four formats: traditional CoT, iCoT, iPoT, or iGraph.
@@ 4.2 Evaluation Metrics
We collect quantitative and qualitative measures to capture both performance outcomes and user experiences: i) Verification Accuracy [nguyen2024interpretable]: For each explanation, we record whether participants correctly labeled the solution as correct or incorrect, providing a direct indicator of how effectively each interface supports error detection and overall comprehension; ii) Error Localization [kamoi2024evaluating]: In cases where the user identifies an explanation as incorrect, they are required to indicate the specific step at which the error occurred. This allows us to assess not only error detection but also participants’ ability to pinpoint the exact reasoning breakdown (avoiding the user answering by guessing); iii) Response Time [nguyen2025hot]: The system automatically logs the total time spent on each explanation, which we use as a proxy for cognitive effort, with longer times potentially indicating greater difficulty in comprehension or in identifying error; iv) Interaction Behavior [ooge2022explaining]: All interface interactions, including play, pause, step forward, and step back, are recorded at the click level. These logs gave us insights into how participants navigate the explanations; and v) Subjective Feedback [kim2023help]: After completing the tasks, participants rate their assigned interface on dimensions such as clarity, ease of use, and overall satisfaction using Likert-scale questions. They are also invited to provide open-ended feedback about what they found helpful or confusing, helping us understand participants’ perceptions of strengths and weaknesses in each format.
@@ 4.3 Post survey Questionnaire
After the experiment, participants will complete 2 questionnaires displayed in Appendix B to assess both their interaction experience and their perceptions of the explanation quality. The questionnaires are designed to address the study’s research questions regarding interpretability, efficiency, and engagement across different explanation formats. All items use a 7-point Likert scale (1 = Strongly Disagree, 7 = Strongly Agree) to capture nuanced user perceptions. Two types of questionnaires are given: Design Feature and a Post-Study. The question sets are inspired by established measures of engagement, usability, and interpretability in human-AI interaction research and are adapted for the design features of our interfaces [rong2023towards, kim2023help, zhou2019mhealth].
@@ 5.1 Quantitative Result
Results in Fig. 5 demonstrate that interactive explanation formats substantially improve participants’ ability to verify reasoning accuracy compared to the traditional CoT. In particular, the iGraph interface achieves the highest verification accuracy ($85.6\%$), followed by the iPoT ($82.5\%$) and iCoT ($80.6\%$) formats, whereas CoT only achieves $73.5\%$. This pattern indicates that interactive explanation formats help users better detect and localize logical faults. These findings align with recent work emphasizing the importance of human-centered explainability, which explains understanding as a process of active answer-seeking rather than passive reception [kim2024human].
Next, we use a Kruskal–Wallis ANOVA to examine whether participants’ ability to identify the specific step containing an error varies across the four explanation formats. The analysis reveals a statistically significant difference among formats ($H(3)=13.12,p=0.004$). Pairwise Mann–Whitney U tests with Bonferroni correction show that iCoT($U=300.8,p=0.045$), iPoT ($U=271.5,p=0.021$), and iGraph ($U=153.0,p<0.001$) interfaces all achieve significantly higher wrong-step identification accuracy compared to the traditional CoT format. Results in Fig. 6 show that IGraph has achieved the highest wrong-step identification accuracy ($85.2\%$), followed by iPoT ($80.1\%$) and iCoT ($79.3\%$), compared to only $66.1\%$ reached by traditional CoT. The higher wrong-step identification accuracy on iGraph further suggests that interactive explanations helped users understand each step more thoroughly. Prior works in explainable artificial intelligence (XAI) argue that unstructured explanations often overwhelm users with text and often make it hard for users to keep track of each reasoning step [lakkaraju2019faithful]. Our findings support these arguments by showing that interactive explanations allow users to accurately identify the origin of an error, rather than simply noticing that one exists. Further, visualizing reasoning steps as interconnected nodes and edges transforms purely textual information into a two-dimensional spatial format, making the logic flow easier to follow and understand (this is further verified by our post survey rating results in Fig. 8).